# Discrete Linear Canonical Transform on Graphs: Uncertainty Principle and Sampling

### v2

@@ 5.1 Simulations
We conducted simulations using the graph signal processing toolbox (GSPBox) [41] for two examples. The first example involves a David sensor network with $N=500$ nodes. We set the bandwidth $|\mathcal{F}|=60$, and the $\mathbf{M}$-bandlimited parameters to $\alpha=0.7,\beta=N/2,\xi=0.5k+0.5,k=1,...,N$. The graph signal is the real part of $\mathbf{x}=\mathrm{real}(\mathbf{f}_{1}+0.5\mathbf{f}_{2}+\mathbf{f}_{3})$ obtained through $\mathbf{M}$-bandlimited operation, where $\mathbf{f}_{i}$ represents the $i$-th column of $\mathbf{O^{-M}}$. The second example involves the vehicular traffic data in Rome. We set the bandwidth $|\mathcal{F}|=50$, and the $\mathbf{M}$-bandlimited parameters to $\alpha=0.9,\beta=N,\xi=0.5k+0.7,k=1,...,N$. The graph signal is the real vehicular traffic data in Rome [42] obtained through $\mathbf{M}$-bandlimited operation. The original graph signals are depicted in Fig. 4.
@@
For both examples, we performed sampling and recovery of the graph signals using four sampling frameworks based on five optimal sampling strategies. All signal recovery formulas utilize Eq. (24) from Theorem 2. The accuracy of the recovery was measured using the NMSE. Fig. 5 illustrates the locations of sampled nodes. The number of sampled nodes, $|\mathcal{S}|$, is 60 for (a), (b), (c), and (d), and 50 for (e), (f), (g), and (h). Fig. 6 displays the signals recovered through Laplacian-based GFT sampling in (a) and (e), GFT sampling in (b) and (f), GFRFT sampling in (c) and (g), and GLCT sampling in (d) and (h). The NMSE results for the David sensor network and Roman vehicular traffic data, across four sampling frameworks employing five optimal strategies, are detailed in Tables 2 and 3, respectively. Note that the parameters selected for our GFRFT and GLCT are consistent with those specified in the aforementioned two examples. Since both A-optimal and T-optimal designs target optimizing the trace of an operator, they yield identical results. Meanwhile, MinPinv and MaxSigMin, essentially embodying E-optimal design, produce similar outcomes. It is evident that Laplacian-based GFT, the GFT, and the GFRFT sampling fail to perfectly recover the graph signals, as these signals, while $\mathbf{M}$-bandlimited, are not strictly bandlimited in the Laplacian-based GFT, the GFT and the GFRFT domains. Only the GLCT sampling achieves perfect signal recovery.
@@
MinFro $6.3526\times 10^{11}$ $4.6075\times 10^{13}$ $4.1884$ $4.5432\times 10^{-22}$
MaxVol $2.6311\times 10^{15}$ $6.0047\times 10^{6}$ $3.2970$ $3.7001\times 10^{-24}$
MinPinv $0.9631$ $1.6822$ $1.5754$ $5.4518\times 10^{-25}$
MaxSigMin $0.9635$ $1.7311$ $1.7956$ $1.8202\times 10^{-25}$
MaxSig $6.3526\times 10^{11}$ $4.6075\times 10^{13}$ $4.1884$ $4.5432\times 10^{-22}$
@@
MinFro $1.2721\times 10^{14}$ $5.1085\times 10^{11}$ $1.3539\times 10^{3}$ $8.7596\times 10^{-23}$
MaxVol $2.9563\times 10^{11}$ $4.1516\times 10^{13}$ $15.4921$ $4.0667\times 10^{-24}$
MinPinv $1.0901$ $1.0038$ $1.0031$ $1.3541\times 10^{-25}$
MaxSigMin $1.0888$ $1.0038$ $1.0086$ $1.7686\times 10^{-25}$
MaxSig $1.2721\times 10^{14}$ $5.1085\times 10^{11}$ $1.3539\times 10^{3}$ $8.7596\times 10^{-23}$
@@ Acknowledgments
This work were supported by the National Natural Science Foundation of China [No. 62171041], the BIT Research and Innovation Promoting Project [No.2023YCXY053], and Natural Science Foundation of Beijing Municipality [No. 4242011]. The authors would like to thank Linyu Peng, Associate Professor in the Department of Mechanical Engineering, Faculty of Science and Technology, Keio University, for his careful guidance.

### v3

@@ 5.1 Simulations
We conducted simulations using the graph signal processing toolbox (GSPBox) [47] for two examples. The first example involves a David sensor network with $N=500$ nodes. We set the bandwidth $|\mathcal{F}|=60$, and the $\mathbf{M}$-bandlimited parameters to $\alpha=0.7,\beta=N/2,\xi=0.5k+0.5,k=1,...,N$. The graph signal $\mathbf{x}$ is the real part of $\mathbf{f}_{1}+0.5\mathbf{f}_{2}+2\mathbf{f}_{3}$, i.e., $\mathbf{x}=\mathrm{real}(\mathbf{f}_{1}+0.5\mathbf{f}_{2}+2\mathbf{f}_{3})$ obtained through $\mathbf{M}$-bandlimited operation, where $\mathbf{f}_{i}$ represents the $i$-th column of $\mathbf{O^{-M}}$. The second example involves the vehicular traffic data in Rome. We set the bandwidth $|\mathcal{F}|=50$, and the $\mathbf{M}$-bandlimited parameters to $\alpha=0.9,\beta=N,\xi=0.5k+0.7,k=1,...,N$. The graph signal is the real vehicular traffic data in Rome [48] obtained through $\mathbf{M}$-bandlimited operation. The original graph signals are depicted in Fig. 6.
@@
For both examples, we performed sampling and recovery of the graph signals using four sampling frameworks based on seven optimal sampling strategies. All signal recovery formulas utilize Eq. (22) from Theorem 2. The accuracy of the recovery was measured using the NMSE. Fig. 7 illustrates the locations of sampled nodes using MinPinv sampling strategy. The number of sampled nodes, $|\mathcal{S}|$, is 60 for (a), (b), (c), and (d), and 50 for (e), (f), (g), and (h). Fig. 8 displays the signals recovered through Laplacian-based GFT sampling in (a) and (e), GFT sampling in (b) and (f), GFRFT sampling in (c) and (g), and GLCT sampling in (d) and (h). The NMSE results for the David sensor network and Roman vehicular traffic data, across four sampling frameworks employing seven optimal strategies, are detailed in Tables 2 and 3, respectively. Note that the parameters selected for our GFRFT and GLCT are consistent with those specified in the aforementioned two examples.
@@
Since both A-optimal and T-optimal designs target optimizing the trace of an operator, MinFro and MaxSig yield identical results. Although MaxVertex is also the A-optimal design, it uses different optimal objective function and has different results. Meanwhile, MaxSpec and MaxVol essentially reflect D-optimal design, yielding similar results. Likewise, MinPinv and MaxSigMin essentially embodying E-optimal design, produce similar outcomes. It is evident that Laplacian-based GFT, the GFT, and the GFRFT sampling fail to perfectly recover the graph signals, as these signals, while $\mathbf{M}$-bandlimited, are not strictly bandlimited in the Laplacian-based GFT, the GFT and the GFRFT domains. Only the GLCT sampling achieves perfect signal recovery.
@@
MinFro $5.5784\times 10^{12}$ $1.6643\times 10^{15}$ $1.7736$ $3.4424\times 10^{-21}$
MaxVertex $3.9576\times 10^{12}$ $9.1507\times 10^{9}$ $471.0998$ $8.0523\times 10^{-19}$
MaxSpec $6.7287\times 10^{12}$ $5.9499\times 10^{12}$ $1.7736$ $1.0796\times 10^{-19}$
MaxVol $6.7287\times 10^{12}$ $1.6188\times 10^{7}$ $5.6375$ $1.2084\times 10^{-22}$
MinPinv $1.0260$ $1.5349$ $1.7736$ $2.9277\times 10^{-24}$
MaxSigMin $1.0024$ $1.4591$ $1.8291$ $3.0375\times 10^{-24}$
MaxSig $5.5784\times 10^{12}$ $1.6643\times 10^{15}$ $1.7736$ $3.4424\times 10^{-21}$
@@
MinFro $1.9346\times 10^{12}$ $2.1664\times 10^{11}$ $5.1485\times 10^{3}$ $1.8502\times 10^{-23}$
MaxVertex $5.1846\times 10^{17}$ $2.2107\times 10^{13}$ $1.1505\times 10^{5}$ $1.3938\times 10^{-22}$
MaxSpec $6.5921\times 10^{11}$ $9.8408\times 10^{12}$ $19.5859$ $3.7200\times 10^{-19}$
MaxVol $6.5921\times 10^{11}$ $9.8408\times 10^{12}$ $20.3122$ $1.7204\times 10^{-23}$
MinPinv $1.0305$ $1.0860$ $1.1172$ $5.2989\times 10^{-25}$
MaxSigMin $1.0281$ $1.0902$ $1.1216$ $1.8111\times 10^{-24}$
MaxSig $1.9346\times 10^{12}$ $2.1664\times 10^{11}$ $5.1485\times 10^{3}$ $1.8502\times 10^{-23}$
@@ Appendix B Proof of Lemma 2
Appendix C Proof of Theorem 3
Let $\mathbf{J=RD}$, we have
thus $\mathbf{J=PD}$ is a projection operator.
Since $\mathbf{RD}$ is a projection operator, $\mathbf{RDx}$ is an approximation of $\mathbf{x}$ in the space of $\mathrm{BL}_{|\mathcal{F}|}\left(\mathbf{O}^{\mathbf{M}}\right)$. We achieve perfect recovery when $\mathbf{x}$ is in the space of $\mathrm{BL}_{|\mathcal{F}|}\left(\mathbf{O}^{\mathbf{M}}\right)$.
@@ Acknowledgments
This work were supported by the National Natural Science Foundation of China [No. 62171041], the BIT Research and Innovation Promoting Project [No.2023YCXY053], and Natural Science Foundation of Beijing Municipality [No. 4242011]. The authors would like to thank the editor, reviewers, and Prof. Linyu Peng of Keio University for their comments that improved the technical quality of this work.